OpenStudy (anonymous):

What is the inverse of the given relation? y = 7x2 – 3

OpenStudy (dumbcow):

basically solve for x and you will get inverse

Parth (parthkohli):

... and then replace \(x\) with \(y\) and vice-versa.

OpenStudy (hba):

f(x)=y=7x^2-3

OpenStudy (anonymous):

You're all talking a foreign language to me. I hate math, more than anything. I don't get this, at all.

OpenStudy (hba):

Hating it and getting it are two different things.

OpenStudy (anonymous):

I hate it, and dont get it.

OpenStudy (dumbcow):

when i said "solve for x" i meant rearrange the terms so that x is by itself \[y = 7x^{2} -3\] add 3 to both sides \[y+3 = 7x^{2}\] ... until x = ?

OpenStudy (anonymous):

7^2 = 49, so 49x?

OpenStudy (dumbcow):

hmm no the 7 is not being squared, only the x

OpenStudy (dumbcow):

let me try this example... 5x = 4 what does x equal?

OpenStudy (anonymous):

wouldn't i divide 5 by both sides?

OpenStudy (dumbcow):

thats right...ok now apply that concept to y+3 = 7x^2

OpenStudy (dumbcow):

or 7x^2 = y+3 if that looks nicer